# Counterflow of lattice polarons in harmonically confined optical lattices

### v1

@@ Appendix B Numerical details
Following Ref. [58], the calculations shown in the main text employ $V_{\text{ho}}=0.008t$ and $N_{b}=40$, resulting in the chosen characteristic density of approximately $\tilde{\rho}_{b}=3.58$. For benchmarking, we also show in this supplemental material results from DMRG calculations for the same density $\tilde{\rho}_{b}$ but with $V_{\text{ho}}=0.0205t$ and $N_{b}=25$. For both parameter choices, including the main text, we consider lattices with $M=N_{b}+20$ sites. These large lattice sizes enable us to avoid any artificial border effect. To aid the convergence, the DMRG results shown in this work also added a very small tilt to the lattice so that, if the profiles develop a finite average value, the impurity prefers to localize at the right, while the bath prefers to localize at the left. However, we have also performed calculations without such tilt, resulting in identical results, but with the impurity (bath) localizing randomly at either the left (right) side or vice versa.
In this supplemental material, and for additional benchmarking, we also show some exact diagonalization (ED) [81] calculations on small lattices with $V_{\text{ho}}=0.2t$, $N_{b}=8$, and $M=14$. While such small lattices show smooth crossovers instead of well-defined transitions, they provide a good crosscheck for our DMRG results, enabling us to rule out any convergence issue in our simulations.
@@ Appendix C Mott insulator–superfluid transition and bath parameters
Figure 5: Results for the one-component Bose-Hubbard model (10) with $\tilde{\rho}=3.58$. (a–c): Density profiles $n$ as a function of $x_{i}$ for the values of $t/U$ indicated in each panel. (d): Central density $n_{c}=n(i=0)$ as a function of $t/U$. Panels (a–d) include ED calculations with $N=8$ and $V_{\text{ho}}=0.2t$, DMRG calculations with $N=25$ and $V_{\text{ho}}=0.0205t$, and DRMG calculations with $N=40$ and $V_{\text{ho}}=0.008t$, as indicated by the legends. (e): One-body correlator as a function of $x_{i}$ for the values of $t/U$ indicated by the legend. Panel (e) only shows DMRG calculations with $N=40$. The dashed black lines are algebraic fits for $t/U=0.3$ and $t/U=0.2$, and the dash-dotted black line is an exponential fit for $t/U=0.3$.
@@
$\mathcal{C}_{\text{SF}}(i)=\langle\hat{a}_{0}\hat{a}^{\dagger}_{i}\rangle.$ (11)
@@ Appendix D Comparison between DRMG and ED calculations
Figure 6: Polaron residue (a,b), polaron energy (c,d), and the average distance between the bosons and the impurity (e,f) as a function of $U_{bI}/U_{bb}$ for $t/U_{bb}=0.2$ (left panels) and $t/U_{bb}=0.1$ (right panels). The panels include ED calculations with $N=8$ and $V_{\text{ho}}=0.2t$, DMRG calculations with $N=25$ and $V_{\text{ho}}=0.0205t$, and DRMG calculations with $N=40$ and $V_{\text{ho}}=0.008t$, as indicated by the legends. In the right panels, the region $0.425\lesssim U_{bI}/U_{bb}\leq 1$ corresponds to the counterflow phase.
@@ Appendix E Impurity-hole model
Here we provide a more detailed presentation of the impurity-hole model. As stated in the main text, we start with an MI state for the bath, $|\text{MI}\rangle=\prod_{i}\hat{a}^{\dagger}_{i,b}|\emptyset\rangle$, in a chain with $M_{\text{CF}}$ sites. An elementary impurity-hole state is then defined by $|i_{\text{IH}}\rangle=\hat{a}^{\dagger}_{i,I}\hat{a}_{i,b}|\text{MI}\rangle$, resulting in the wavefunction $|\Psi_{\text{CF}}\rangle=\sum_{i}\alpha_{i}|i_{\text{IH}}\rangle$, where $\sum_{i}\alpha^{2}_{i}=1$. Note that this wavefunction does not conserve the number of bosons, and thus is only an approximation for a large number of particles. Naturally, this restriction stems from the consideration of a chain in this simple model. In the full model, the addition of the harmonic trap provides the additional site needed to create the annihilated boson.
@@
Finally, the correlator reads
$\mathcal{C}_{\text{SF}}(i)=$ $\langle\Psi_{\text{CF}}|\hat{a}_{0,b}\hat{a}^{\dagger}_{0,I}\hat{a}^{\dagger}_{i,b}\hat{a}_{i,I}|\Psi_{\text{CF}}\rangle$
@@
Then, by imposing $n_{I}(i)=\alpha_{i}^{2}=n_{I}^{(0)}\cos^{2}(i\pi/M_{\text{CF}})$, one gets $\mathcal{C}_{\text{SCF}}(i)=n_{I}^{(0)}\cos(i\pi/M_{\text{CF}})$. Here $M_{\text{CF}}=2/n_{I}^{(0)}$, so that there is only one impurity in the system ($\sum_{i}n_{I}(i)=1$). Note that this sum of the profile needs to be performed in $i$ instead of the rescaled length $x_{i}$, explaining the value $\ell_{\text{CF}}=2/(\xi n_{I}^{(0)})$ reported in the main text.

### v2

@@ Appendix B Numerical details
Following Ref. [58], the calculations shown in the main text employ $V_{\text{ho}}=0.008t$ and $N_{b}=40$, resulting in the chosen characteristic density of approximately $\tilde{\rho}_{b}=3.58$. For benchmarking, we also show in this supplemental material results from DMRG calculations for the same density $\tilde{\rho}_{b}$ but with $V_{\text{ho}}=0.0205t$ and $N_{b}=25$. For both parameter choices, including the main text, we consider lattices with $M=N_{b}+20$ sites. These large lattice sizes enable us to avoid any artificial border effect. To aid the convergence, the DMRG results shown in this work also added a very small tilt to the lattice so that, if the profiles develop a finite average value, the impurity prefers to localize at the right, while the bath prefers to localize at the left. However, we have also performed calculations without such a tilt, resulting in identical results, but with the impurity (bath) localizing randomly at either the left (right) side or vice versa.
In this supplemental material, and for additional benchmarking, we also show some exact diagonalization (ED) [82] calculations on small lattices with $V_{\text{ho}}=0.2t$, $N_{b}=8$, and $M=14$. While such small lattices show smooth crossovers instead of well-defined transitions, they provide a good cross-check for our DMRG results, enabling us to rule out any convergence issue in our simulations.
@@ Appendix C Mott insulator–superfluid transition and bath parameters
Figure 5: Results for the one-component Bose-Hubbard model (10) with $\tilde{\rho}=3.58$. (a–c): Density profiles $n$ as a function of $x_{i}$ for the values of $t/U$ indicated in each panel. (d): Central density $n_{c}=n(i=0)$ as a function of $t/U$. Panels (a–d) include ED calculations with $N=8$ and $V_{\text{ho}}=0.2t$, DMRG calculations with $N=25$ and $V_{\text{ho}}=0.0205t$, and DRMG calculations with $N=40$ and $V_{\text{ho}}=0.008t$, as indicated in the legends. (e): One-body correlator as a function of $x_{i}$ for the values of $t/U$ indicated in the legend. Panel (e) only shows DMRG calculations with $N=40$. The dashed black lines are algebraic fits for $t/U=0.3$ and $t/U=0.2$, and the dash-dotted black line is an exponential fit for $t/U=0.3$.
@@
$\mathcal{C}^{(b)}_{\text{SF}}(i)=\langle\hat{a}_{0}\hat{a}^{\dagger}_{i}\rangle.$ (11)
@@ Appendix D Comparison between DRMG and ED calculations
Figure 6: Polaron residue (a,b), polaron energy (c,d), and the average distance between the bosons and the impurity (e,f) as a function of $U_{bI}/U_{bb}$ for $t/U_{bb}=0.2$ (left panels) and $t/U_{bb}=0.1$ (right panels). The panels include ED calculations with $N=8$ and $V_{\text{ho}}=0.2t$, DMRG calculations with $N=25$ and $V_{\text{ho}}=0.0205t$, and DRMG calculations with $N=40$ and $V_{\text{ho}}=0.008t$, as indicated in the legends. In the right panels, the region $0.425\lesssim U_{bI}/U_{bb}\leq 1$ corresponds to the counterflow phase.
@@ Appendix E Impurity-hole model
Here we provide a more detailed presentation of the impurity-hole model. As stated in the main text, we start with an MI state for the bath, $|\text{MI}\rangle=\prod_{i}\hat{a}^{\dagger}_{i,b}|\emptyset\rangle$, in a chain with $M_{\text{CF}}$ sites. An elementary impurity-hole state is then defined by $|i_{\text{IH}}\rangle=\hat{a}^{\dagger}_{i,I}\hat{a}_{i,b}|\text{MI}\rangle$, resulting in the wavefunction $|\Psi_{\text{CF}}\rangle=\sum_{i}\alpha_{i}|i_{\text{IH}}\rangle$, where $\sum_{i}|\alpha_{i}|^{2}=1$. Note that this wavefunction does not conserve the number of bosons, and thus is only an approximation for a large number of particles. Naturally, this restriction stems from the consideration of a chain in this simple model. In the full model, the addition of the harmonic trap provides the additional site needed to create the annihilated boson.
@@
Finally, the anti-pair correlator reads
$\mathcal{C}_{\text{AP}}(i)=$ $\langle\Psi_{\text{CF}}|\hat{a}_{0,b}\hat{a}^{\dagger}_{0,I}\hat{a}^{\dagger}_{i,b}\hat{a}_{i,I}|\Psi_{\text{CF}}\rangle$
@@
By now imposing that the coefficients $\alpha_{i}$ are real and positive, from $n_{I}(i)=|\alpha_{i}|^{2}=n_{I}^{(0)}\cos^{2}(i\pi/M_{\text{CF}})$ we get that
$\alpha_{i}=\sqrt{n_{I}^{(0)}}\cos(i\pi/M_{\text{CF}}).$ (19)
With this, the correlator finally takes the form $\mathcal{C}_{\text{AP}}(i)=n_{I}^{(0)}\cos(i\pi/M_{\text{CF}})$. Here $M_{\text{CF}}=2/n_{I}^{(0)}$, so that there is only one impurity in the system ($\sum_{i}n_{I}(i)=1$). Note that this sum of the profile needs to be performed in $i$ instead of the rescaled length $x_{i}$, explaining the value $\ell_{\text{CF}}=2/(\xi n_{I}^{(0)})$ reported in the main text.